OpenStudy (anonymous):

Solve: 4x + 16 + 40 = 160

OpenStudy (anonymous):

4x = 120 - 16 x = 30 -4 x=26

OpenStudy (anonymous):

104/4=26

OpenStudy (anonymous):

\[4x+56=160\] \[4x=104\] \[x=26\]